### v1

# A Stochastic Genetic Interacting Particle Method for Reaction-Diffusion-Advection Equations

Abstract
We develop and analyze a stochastic genetic interacting particle method (SGIP) for reaction-diffusion-advection (RDA) equations. The method employs operator splitting to approximate the advection-diffusion and reaction processes, treating the former by particle drift-diffusion and the latter by exact or implicit integration of reaction dynamics over bins where particle density is estimated by a histogram. A key innovation is the incorporation of adaptive resampling to close the loop of particle and density field description of solutions, mimicking mutation in genetics. Resampling is also critical for maintaining long-term stability by redistributing particles with the evolving density field. We provide a comprehensive error analysis, and establish rigorous convergence bounds under appropriate regularity assumptions. Numerical experiments from one to three space dimensions demonstrate the method’s effectiveness across various reaction types (FKPP, cubic, Arrhenius) and flow configurations (shear, cellular, cat’s eye, ABC flows), showing excellent agreement with finite difference method (FDM) while offering computational advantages for complex flow geometries and higher-dimensional problems.
Keywords: Stochastic particle method, resampling, reaction-diffusion-advection equations, operator splitting, convergence analysis, 3D simulations.
1 Introduction
Reaction-Diffusion-Advection (RDA) equations play a fundamental role in modeling numerous physical, biological, and chemical processes. The general form of RDA equations is given by:
$\frac{\partial u}{\partial t}+\nabla\cdot(\mathbf{v}u)=D\Delta u+r(u)$ (1)
where $u(\mathbf{x},t)$ represents the concentration or density field, $\mathbf{v}(\mathbf{x},t)$ is ambient fluid flow field, $D$ is molecular diffusion coefficient, and $r(u)$ is reaction nonlinearity. The versatility of RDA equations is reflected in their wide-ranging applications [6, 24, 35]. In ecology, the Fisher-Kolmogorov-Petrovsky-Piskunov (FKPP) equation models propagating population and gene waves [11], while developmental biology employs Turing’s reaction-diffusion model to explain morphological patterning [33]. The framework also extends to materials science for simulating phase separation via the Cahn-Hilliard equation [4], and to finance, where it underpins the Black-Scholes model for option pricing [3], collectively underscoring its broad utility [18].
The development of efficient numerical methods for these equations remains an active research area, with challenges from nonlinearities, multiscale phenomena, and high-dimensionality. The Finite Difference Method (FDM) is well-suited for simple geometries, with established stability and convergence properties [21, 14]. For complex domains, the Finite Element Method (FEM) offers flexibility, grounded in variational principles [20, 17, 28]. Operator splitting techniques (e.g., Lie–Trotter and Strang splitting) are often used to mitigate stiffness from diffusion or reaction terms [32, 31]. Despite their successes, the mesh-based methods struggle with the curse of dimensionality and problems involving large deformations or moving boundaries where mesh maintenance becomes prohibitively expensive.
Particle (Lagrangian) techniques present a mesh-free alternative to handle large deformation (e.g. sharp internal layers) and high-dimensional problems. In the 1980’s, [29] [29, 30] introduced a random walk method on the spatial derivative of solutions. The reaction is treated by a reproduction and destruction process to evolve steep wavefronts with an application to Hodgkin–Huxley equations. Representing spatial derivative by an empirical measure of particles, Puckett [26] showed its convergence for the FKPP equation. These early works are limited to one space (1D) dimension and monotone initial data however. For recent efforts on particle methods of RDA type equations, we refer to [22, 19, 37, 23] and references therein. In general, a particle description alone can be either computationally expensive or insufficient to represent solutions of physical and biological systems. In the absence of low order reactions, the current authors [34, 15, 16] recently introduced efficient stochastic particle-field methods to compute parabolic-parabolic chemotaxis and haptotaxis advection-diffusion systems up to 3D, in lieu of a history dependent pure particle description involving heat kernels.
In the spirit of particle-field representation, the stochastic genetic interacting particle (SGIP) method developed in this work provides an efficient and robust framework for RDA equations in several space dimensions. It is a natural extension of the SGIP [22, 37] built from the Feynman-Kac semigroup for linear RDA equations. The algorithm consists of a 3-step operator-splitting to decompose the RDA process: drift-diffusion, density update and particle resampling. The advection-diffusion is handled by drift-diffusion (Ito equation) of particles. The reaction process is approximated by going between particles and density: the density field is a histogram counting particle positions over bins, and is evolved in time via reaction ODEs by either an exact (e.g. for FKPP) or a stable implicit integration (Algorithm 2). Then particles are redrawn from the updated density by an adaptive resampling step (Algorithm 3) so that they are ready for advection-diffusion again to close the loop. The joint particle-density evolution is a way to overcome the difficulty of seeking a self-contained particle description. In case of SGIP for linear RDA [22, 37], the particle dynamics is in a close loop via drift-diffusion and resampling steps without relying on density field, due to reaction rate being a known potential function. The resampling step here mimics mutation in genetics in that particles are more likely to survive in a bin to the next generation if the bin has more particles (is more healthy) than its neighbors. Resampling is also critical for long-term stability and computational efficiency by maintaining a well-balanced representation of the evolving solution, i.e. keeping variance in check over time. Our SGIP framework here is supported by a rigorous convergence analysis under appropriate regularity assumptions, which establishes consistent accuracy controlled by time step, spatial resolution, and the number of particles.
Our numerical experiments comprehensively demonstrate SGIP’s capabilities across various dimensions and complexities. In 1D scenarios, SGIP shows excellent agreement with FDM for different reaction types (FKPP, cubic, and Arrhenius) [10, 1], accurately capturing propagation speeds and front shapes. The method effectively resolves a spectrum of complex flow configurations, from fundamental 2D shear, cellular and cat’s eye flows to challenging 3D Arnold-Beltrami-Childress (ABC) flows [27, 7, 25, 9, 36], even when coupled with Lipschitz-continuous reaction terms. In these scenarios, particularly in 3D, SGIP maintains high accuracy while offering significant computational advantages over traditional mesh-based methods like the FDM. As the diffusion coefficient decreases and advection dominates, the SGIP framework remains stable and efficient, in stark contrast to FDM, which suffers from prohibitive computational costs and severe numerical instability under such conditions.
The SGIP method provides a simple-to-program, scalable, low-cost and stable framework for high-dimensional RDA problems. Its mesh-free formulation bypasses the geometric constraints of grid-based methods, allowing it to handle complex flows and large gradients of solutions with ease. This capability is augmented by favorable computational scaling in high dimensions and long-term stability, thanks to adaptive resampling. These combined strengths allow SGIP to open new avenues for simulating complex multiscale phenomena in physical and biological systems, particularly in scenarios involving intricate flow geometries and requiring long-time integration.
The rest of the paper is organized as follows. In Section 2, we present the detailed formulation of the SGIP algorithm. This section is divided into three subsections dedicated to the stochastic advection-diffusion step, the reaction step, and the adaptive resampling strategy, respectively. Section 3 is devoted to the theoretical analysis, where a comprehensive error decomposition is provided and convergence bounds are rigorously established. In Section 4, extensive numerical experiments are conducted to validate the method’s performance, ranging from 1D benchmark tests to 2D and 3D simulations involving complex flows. Finally, we conclude with a summary of our findings and a discussion of future research directions in Section 5.
2 SGIP Algorithm
In this section, we will present the SGIP algorithm for solving the reaction-diffusion-advection (RDA) equations. The mathematical foundation of our method is the operator-splitting technique. We consider the RDA equation as an abstract evolution equation:
$\frac{\partial u}{\partial t}=(\mathcal{A}+\mathcal{R})u,\quad u(0)=u_{0}$ (2)
where $\mathcal{A}u=\nabla\cdot(\mathbf{v}u)+D\Delta u$ is the advection-diffusion operator and $\mathcal{R}u=r(u)$ is the reaction operator. Following the Lie-Trotter splitting scheme, we approximate the solution evolution by decoupling the operators:
$u_{n+1}\approx\mathcal{S}_{\mathcal{R}}^{\Delta t}\circ\mathcal{S}_{\mathcal{A}}^{\Delta t}(u_{n})$ (3)
Here, $u_{n}$ denotes the numerical approximation of $u(\mathbf{x},t_{n})$ at time $t_{n}$, and $u_{n+1}$ denotes the approximation at the next time level $t_{n+1}=t_{n}+\Delta t$. This splitting enables us to employ specialized numerical techniques for each physical process, ensuring its long-term stability and efficiency.
For the advection-diffusion part $\mathcal{S}_{\mathcal{A}}^{\Delta t}$, we employ the stochastic particle method, which turns out to be stable in various high-dimensional advection-dominant problems. For the reaction $\mathcal{S}_{\mathcal{A}}^{\Delta t}$ counterpart, we apply a density-based approach, which recovers the density $u$ from the particle formulation and approximates the reaction $\mathcal{R}$ by resampling. Details of each part of the algorithm are listed in the following subsections.
2.1 Initialization
Let $\Omega=[-L,L]^{d}$ be the computational domain. We approximate the density field $u(\mathbf{x},t)$ using an empirical measure of $N$ particles:
$u(\mathbf{x},t)\approx\frac{M_{0}}{N}\sum_{i=1}^{N}\delta(\mathbf{x}-\mathbf{X}^{i}(t))$ (4)
where $M_{0}=\int_{\Omega}u_{0}(\mathbf{x})d\mathbf{x}$ is the initial total mass and $\{\mathbf{X}^{i}(t)\}_{i=1}^{N}$ are particle positions. The initial particle positions at $t_{0}=0$ are sampled from the normalized initial distribution:
$\mathbf{X}_{0}^{i}\sim\rho_{0}(\mathbf{x})=\frac{u_{0}(\mathbf{x})}{M_{0}},\quad i=1,\dots,N$ (5)
Particle masses are initialized uniformly according to:
$m_{0}^{i}=\frac{M_{0}}{N},\quad i=1,\dots,N.$ (6)
In the temporal direction, we assume that the computational time range $[0,T]$ is partitioned into time steps by $\{t_{n}\}_{n=0}^{n_{T}}$, where $t_{0}=0$ and $t_{n_{T}}=T$. We consider the uniform time step, i.e., there exists $\Delta t=t_{n+1}-t_{n}$ for $n=0,\dots,n_{T}-1$. For ease of presenting our algorithm, we will use a slight abuse of notation. We will represent the density $u$ at time $t_{n}$ as $u_{n}$, and the particle positions at time $t_{n}$ as $\{\mathbf{X}_{n}^{i}\}_{i=1}^{N}$. Similarly, particle masses at time $t_{n}$ are denoted as $m_{n}^{i}$.
2.2 Advection-Diffusion Step
The advection-diffusion sub-step $\mathcal{S}_{\mathcal{A}}^{\Delta t}$ advances the solution from $t_{n}$ to an intermediate time. Each particle follows the stochastic differential equation:
$d\mathbf{X}^{i}=\mathbf{v}(\mathbf{X}^{i},t)dt+\sqrt{2D}d\mathbf{W}^{i}(t)$ (7)
where $\mathbf{v}(\mathbf{x},t)$ is the velocity field and $\mathbf{W}^{i}(t)$ are independent Brownian motions.
Discretized using the Euler-Maruyama scheme over the time interval $t_{n},t_{n+1}]$:
$\mathbf{X}_{*}^{i}=\mathbf{X}_{n}^{i}+\mathbf{v}(\mathbf{X}_{n}^{i},t_{n})\Delta t+\sqrt{2D\Delta t}\boldsymbol{\xi}_{n}^{i}$ (8)
where $\boldsymbol{\xi}_{n}^{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$ are independent standard normal random variables, and $\mathbf{X}_{*}^{i}$ denotes the intermediate particle positions.
Algorithm 1 Advection-Diffusion Update
1:Particle positions $\{\mathbf{X}_{n}^{i}\}_{i=1}^{N}$, velocity field $\mathbf{v}(\mathbf{x},t)$, time step $\Delta t$
2:Intermediate particle positions $\{\mathbf{X}_{*}^{i}\}_{i=1}^{N}$
3:for $i=1$ to NN do
4:  Sample $\boldsymbol{\xi}_{n}^{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$
5:  $\mathbf{X}_{*}^{i}\leftarrow\mathbf{X}_{n}^{i}+\mathbf{v}(\mathbf{X}_{n}^{i},t_{n})\Delta t+\sqrt{2D\Delta t}\boldsymbol{\xi}_{n}^{i}$
2.3 Reaction Step
The reaction sub-step $\mathcal{S}_{\mathcal{R}}^{\Delta t}$ handles the nonlinear reaction term from the intermediate state to time $t_{n+1}$. This step consists of two main components: density estimation on a Eulerian grid, followed by reaction integration.
2.3.1 Density Estimation
To evaluate the reaction terms, we first reconstruct the density field from the particle distribution. The domain $\Omega=[-L,L]^{d}$ is discretized into $M^{d}$ uniform bins with centers $\{\mathbf{x}_{j}\}_{j=1}^{M^{d}}$ and bin size $\Delta x=2L/M$. The density in each bin is estimated by counting particles and summing their masses:
$\hat{u}_{*}^{j}=\frac{1}{(\Delta x)^{d}}\sum_{i:\mathbf{X}_{*}^{i}\in\text{bin}_{j}}m_{n}^{i}$ (9)
This provides a piecewise constant approximation of the density field at the intermediate state, which serves as the initial condition for the reaction integration.
2.3.2 Reaction Integration
Given the density estimates $\{\hat{u}_{*}^{j}\}$, we solve the reaction equation $\partial_{t}u=r(u)$ over the time interval $\Delta t$ independently in each bin. We consider two cases based on the form of the reaction term:
Closed-form Solutions
For certain reaction terms, exact integration is possible. A prominent example is the FKPP reaction:
r(u)=u(1u)r(u)=u(1-u) (10)
which admits the exact solution:
$\hat{u}_{n+1}^{j}=\frac{\hat{u}_{*}^{j}e^{\Delta t}}{1+\hat{u}_{*}^{j}(e^{\Delta t}-1)}$ (11)
Other examples include linear reactions $r(u)=\lambda u$, which yield exponential growth/decay $\hat{u}_{n+1}^{j}=\hat{u}_{*}^{j}e^{\lambda\Delta t}$.
General Reaction Terms
For reaction terms without closed-form solutions, such as:
  • Cubic reaction: $r(u)=u^{2}(1-u)$,
  • Arrhenius-type reaction: $r(u)=e^{-E/u}(1-u)$ with activation energy $E>0$,
  • Polynomial reactions: $r(u)=\sum_{k=0}^{K}a_{k}u^{k}$,
we employ implicit time integration methods to maintain numerical stability. The backward Euler scheme reads:
$\frac{\hat{u}_{n+1}^{j}-\hat{u}_{*}^{j}}{\Delta t}=r(\hat{u}_{n+1}^{j})$ (12)
which is solved numerically using Newton’s method. For stiff problems, higher-order implicit methods such as Crank-Nicolson [8] or BDF2 [12] can be employed.
The reaction integration is performed independently in each spatial bin, making this step highly parallelizable. After reaction integration, we obtain the updated density field $\{\hat{u}_{n+1}^{j}\}$ at time $t_{n+1}$.
Algorithm 2 Density Estimation and Reaction Integration
1:Intermediate particles $\{\mathbf{X}_{*}^{i}\}$, masses $\{m_{n}^{i}\}$, reaction function $r(u)$, $\Delta t$
2:Updated density field $\{\hat{u}_{n+1}^{j}\}_{j=1}^{M^{d}}$, masses $\{m_{n+1}^{i}\}$
3:Density Estimation:
4:for each bin $j$ do
5:  $\hat{u}_{*}^{j}\leftarrow\frac{1}{(\Delta x)^{d}}\sum_{i:\mathbf{X}_{*}^{i}\in\text{bin}_{j}}m_{n}^{i}$
6:Reaction Integration:
7:for each bin $j$ do
8:  if $r(u)$ has closed-form solution then
9:   $\hat{u}_{n+1}^{j}\leftarrow\text{ExactIntegration}(\hat{u}_{*}^{j},r,\Delta t)$
10:  else
11:   $\hat{u}_{n+1}^{j}\leftarrow\text{ImplicitIntegration}(\hat{u}_{*}^{j},r,\Delta t)$
12:Mass Update:
[…]
2.4 Resampling Step
Following the reaction integration that produces the updated density field $\{\hat{u}_{n+1}^{j}\}$, we employ a resampling strategy to redistribute particles according to this new density distribution. This step ensures particle efficiency and adaptivity while maintaining mass conservation.
The probability for a particle to be resampled in bin $j$ is proportional to the mass contained in that bin:
$p_{j}=\frac{\hat{u}_{n+1}^{j}\cdot(\Delta x)^{d}}{M_{n+1}},\quad j=1,\dots,M^{d}$ (13)
where $M_{n+1}=\sum_{j=1}^{M^{d}}\hat{u}_{n+1}^{j}\cdot(\Delta x)^{d}$ is the total mass after reaction computed in Alg.2. These probabilities guide the redistribution of particles during the resampling step to match the updated density distribution.
New particle counts are generated from a multinomial distribution:
$(n_{1},\dots,n_{M^{d}})\sim\text{Multinomial}(N,p_{1},\dots,p_{M^{d}})$ (14)
The resampling process considers the current particle distribution to minimize statistical errors:
Algorithm 3 Resampling Procedure
1:Intermediate particle positions $\{\mathbf{X}_{*}^{i}\}_{i=1}^{N}$, updated density field $\{\hat{u}_{n+1}^{j}\}_{j=1}^{M^{d}}$
2:Resampled particles $\{\mathbf{X}_{n+1}^{i}\}_{i=1}^{N}$ with updated masses $\{m_{n+1}^{i}\}_{i=1}^{N}$
[…]
4:Compute $p_{j}=\hat{u}_{n+1}^{j}\cdot(\Delta x)^{d}/M_{n+1}$ for $j=1,\dots,M^{d}$
5:Sample $(n_{1},\dots,n_{M^{d}})\sim\text{Multinomial}(N,p_{1},\dots,p_{M^{d}})$
6:Initialize empty list $\mathbf{X}_{\text{new}}$
7:for $j=1$ to $M^{d}$ do
8:  $c_{j}\leftarrow\left|\{\mathbf{X}_{*}^{i}\in\text{bin}_{j}\}\right|$ $\triangleright$ Current particle count in bin $j$
9:  if $c_{j}>0$ then
10:   $\mathcal{X}_{j}\leftarrow\{\mathbf{X}_{*}^{i}\mid\mathbf{X}_{*}^{i}\in\text{bin}_{j}\}$ $\triangleright$ Get particles in bin $j$
11:   if $n_{j}\leq c_{j}$ then
12:     $\mathbf{X}_{\text{new},j}\leftarrow\text{RandomChoice}(\mathcal{X}_{j},n_{j},\text{without replacement})$
13:   else
14:     $\mathbf{X}_{\text{new},j}\leftarrow\text{RandomChoice}(\mathcal{X}_{j},n_{j},\text{with replacement})$
15:  else
16:   $\mathbf{X}_{\text{new},j}\leftarrow\text{UniformSample}(\text{bin}_{j},n_{j})$ $\triangleright$ Sample uniformly from empty bin
17:  Append $\mathbf{X}_{\text{new},j}$ to $\mathbf{X}_{\text{new}}$
[…]
19:Set $m_{n+1}^{i}=M_{n+1}/N$ for all $i=1,\dots,N$
The resampling strategy employs different sampling methods based on the relationship between target and current particle counts:
  • When $n_{j}\leq c_{j}$: Sample without replacement to preserve diversity
  • When $n_{j}>c_{j}$: Sample with replacement to meet the target count
  • When $c_{j}=0$: Generate new particles uniformly within the bin
This approach minimizes statistical errors and maintains good representation of the density field while adapting to the evolving solution structure.
Combining all components developed in the previous subsections, we present the complete SGIP algorithm for RDA equations in Algorithm 4.
Algorithm 4 SGIP Method for RDA Equations
1:NN, $T$, $\Delta t$, $L$, $d$, $u_{0}(\mathbf{x})$, $\mathbf{v}(\mathbf{x},t)$, $D$, $r(u)$, $M$
2:Estimated density $\{\hat{u}_{n}\}_{n=0}^{n_{T}}$
3:Initialization:
4:Sample $\mathbf{X}_{0}^{i}\sim u_{0}(\mathbf{x})/M_{0}$ for $i=1,\dots,N$
5:Set $m_{0}^{i}=M_{0}/N$ for $i=1,\dots,N$
[…]
7:for $n=0$ to $n_{T}-1$ do
8:  Advection-Diffusion Step:
[…]
10:  Reaction Step:
[…]
12:  Resampling Step:
[…]
14:return $\{\hat{u}_{n}\}_{n=0}^{n_{T}}$
Each time step requires $\mathcal{O}(N)$ operations for particle motion, $\mathcal{O}(N)$ for bin assignment, and $\mathcal{O}(M^{d})$ for density estimation and resampling. The overall complexity per time step is $\mathcal{O}(N+M^{d})$, making the method efficient for moderate-dimensional problems.
Remark 1 (Fast adaptive algorithm for computational efficiency).
We note that for reaction-diffusion equations with FKPP type nonlinearities where the solution remains bounded ($u_{\text{max}}=1$), significant computational savings can be achieved. In future work, we will develop an adaptive acceleration technique that exploits the known propagation behavior of the equation. Specifically, regions where $u=1$ (saturated regions) can be identified and excluded from particle-based computation, particles are then only employed in the non-saturated transition regions where $u<1$. This adaptive approach dramatically reduces the number of active bins $M^{d}_{\text{active}}\ll M^{d}$ required for density estimation and resampling, leading to substantial speedup, particularly in higher-dimensional problems where the curse of dimensionality would otherwise limit practical computations.
3 Error Analysis of the SGIP method for RDA Equation
In this section, we present a rigorous error analysis for the SGIP algorithm applied to the RDA equation. The analysis systematically decomposes the total error into four fundamental components: splitting error, particle approximation error, spatial discretization error, and sampling error. We establish precise bounds for each component and combine them to obtain the overall error estimate.
Let $u(t)$ denote the exact solution of the FKPP equation, $u^{\mathrm{Lie}}(t)$ denote the exact Lie splitting solution (Eqs.(2)(3)), and $\widetilde{u}_{n}$ denote the numerical SGIP solution at time $t_{n}=n\Delta t$.
We analyze the local error of the SGIP by decomposing it into distinct components corresponding to different stages of the algorithm. The one-step error at time $t_{n+1}$ admits the following decomposition:
$\begin{aligned} \|\tilde{u}_{n+1}-u_{n+1}\|_{L^{2}}&\leq\underbrace{\|\tilde{u}_{n+1}-\tilde{u}_{n+1}^{\text{Lie}}\|_{L^{2}}}_{\varepsilon_{\text{num}}}+\underbrace{\|\tilde{u}_{n+1}^{\text{Lie}}-u_{n+1}^{\text{Lie}}\|_{L^{2}}}_{\varepsilon_{\text{propag}}}+\underbrace{\|u_{n+1}^{\text{Lie}}-u_{n+1}\|_{L^{2}}}_{\varepsilon_{\text{split}}}\end{aligned}$ (15)
where $u^{\mathrm{Lie}}_{n+1}=S_{\mathcal{R}}(\Delta t)S_{\mathcal{A}}(\Delta t)u_{n}$ is the Lie splitting solution from the previous exact solution $u_{n}$. Here, $\varepsilon_{\mathrm{split}}$ represents the splitting error due to operator splitting, while $\varepsilon_{\mathrm{num}}$ captures the numerical implementation error arising from the particle discretization, and $\varepsilon_{\mathrm{propag}}$ denotes the error from propagating the numerical solution over one time step using the splitting scheme.
The numerical implementation error $\varepsilon_{\mathrm{num}}$ can be further decomposed by tracing through the algorithmic stages:
$\varepsilon_{\mathrm{num}}$ $=\|\tilde{u}_{n+1}-\widetilde{u}^{\mathrm{Lie}}_{n+1}\|_{L^{2}}$
$\leq\underbrace{\|S_{\mathcal{R}}(\Delta t)(\tilde{u}^{\mathrm{D}}_{n+1})-S_{\mathcal{R}}(\Delta t)S_{\mathcal{A}}(\Delta t)\tilde{u}_{n}\|_{L^{2}}}_{\varepsilon_{\mathrm{dens}}}$
$\quad+\underbrace{\|\tilde{u}^{\mathrm{R}}_{n+1}-S_{\mathcal{R}}(\Delta t)(\tilde{u}^{\mathrm{D}}_{n+1})\|_{L^{2}}}_{\varepsilon_{\mathrm{react}}}$
$\quad+\underbrace{\|\tilde{u}_{n+1}-\tilde{u}^{\mathrm{R}}_{n+1}\|_{L^{2}}}_{\varepsilon_{\mathrm{resamp}}}$ (16)
where the intermediate quantities are:
  • $\widetilde{u}^{\mathrm{D}}_{n+1}$: numerical density after particle advection-diffusion and binning,
  • $\widetilde{u}^{\mathrm{R}}_{n+1}$: numerical density after reaction integration step,
  • $\widetilde{u}_{n+1}$: final numerical solution after resampling step.
The numerical implementation error comprises four distinct components: $\varepsilon_{\mathrm{dens}}$ represents the density estimation error from particle advection-diffusion and spatial binning; $\varepsilon_{\mathrm{react}}$ corresponds to the reaction integration error from numerically solving the reaction ODEs; and $\varepsilon_{\mathrm{resamp}}$ captures the resampling error introduced during particle redistribution.
To establish rigorous error bounds, we introduce the following technical assumptions:
Assumption 1 (Regularity and boundedness).
The solution and operators satisfy:
  1. 1.
    Initial regularity: The initial data $u_{0}$ belongs to the Sobolev space $H^{s}(\Omega)$ with $s>\frac{d}{2}+3$, ensuring sufficient spatial regularity for the error analysis.
  2. 2.
    Velocity field regularity: The velocity field $\mathbf{v}(\mathbf{x},t)$ is uniformly Lipschitz continuous in space:
    $|\mathbf{v}(\mathbf{x},t)-\mathbf{v}(\mathbf{y},t)|\leq L_{v}|\mathbf{x}-\mathbf{y}|\quad\text{for all }\mathbf{x},\mathbf{y}\in\Omega,t\in[0,T],$ (17)
    and is uniformly bounded: $\|\mathbf{v}\|_{L^{\infty}(\Omega\times[0,T])}\leq M_{1}$.
  3. 3.
    Boundedness: The exact solution $u(t,\cdot)$ remains uniformly bounded in the interval $[0,1]$ for all $t\in[0,T]$, and maintains smoothness throughout the time evolution.
  4. 4.
    Semigroup contractivity: The boundary conditions (Neumann, periodic, or decay conditions on $\mathbb{R}^{d}$) are such that the advection-diffusion semigroup $S_{\mathcal{A}}(t)=e^{t\mathcal{A}}$ is contractive in $L^{2}(\Omega)$, i.e.,
    $\|S_{\mathcal{A}}(t)u\|_{L^{2}}\leq\|u\|_{L^{2}}\quad\text{for all }t\geq 0.$ (18)
  5. 5.
    Reaction term regularity: The reaction function $r(u)$ satisfies the Lipschitz condition
    $|r(u)-r(v)|\leq L_{r}|u-v|\quad\text{for all }u,v\in[0,u_{\max}],$ (19)
    and is twice continuously differentiable with bounded second derivative $|r^{\prime\prime}(u)|\leq M_{2}$, ensuring well-posedness of the reaction steps.
Assumption 2 (CFL-like Condition).
In the SGIP algorithm, we assume that the bin size $\Delta x$ approaches 0 faster than the time step $\Delta t$ approaches 0. Additionally, the number of particles $N$ is assumed to grow sufficiently fast relative to $\Delta x$ and $\Delta t$.
More formally, as $\Delta t\to 0$, $\Delta x\to 0$, and $N\to\infty$, there exist parameters $\kappa,\nu$ such that:
$\kappa:=\frac{\Delta x}{\Delta t},\ \kappa\to 0,\qquad\nu:=\frac{1}{\sqrt{N\Delta x^{d}}\Delta t},\ \nu\to 0.$ (20)
We first present the main convergence theorem, with its proof relying on the technical lemmas established in the subsequent subsection.
Theorem 3 (Global error bound).
Under Assumptions 1 and 2, for $n\Delta t\leq T$, the global error of the SGIP algorithm satisfies:
$\sup_{0\leq n\leq T/\Delta t}\mathbb{E}\,\|\widetilde{u}_{n}-u_{n}\|_{L^{2}}\leq C_{T}\left(\Delta t+(N\Delta x^{d}\Delta t)^{-1/2}+\frac{\Delta x}{\Delta t}\right)$ (21)
[…]
The proof follows from the subsequent lemmas that bound each error component.
Lemma 1 (Propagation error estimate).
Under Assumption 1, the propagation error satisfies:
$\|S_{\mathcal{R}}(\Delta t)S_{\mathcal{A}}(\Delta t)\tilde{u}_{n}-S_{\mathcal{R}}(\Delta t)S_{\mathcal{A}}(\Delta t)u_{n}\|_{L^{2}}\leq(1+C_{1}\Delta t)\|\tilde{u}_{n}-u_{n}\|_{L^{2}}$ (22)
where $C_{1}$ depends on the Lipschitz constants of the advection-diffusion and reaction operators.
Proof.
We analyze the propagation error by considering each operator separately.
For the advection-diffusion operator $\mathcal{A}u=-\nabla\cdot(\mathbf{v}u)+D\Delta u$, consider two solutions $u(t)=S_{\mathcal{A}}(t)u_{0}$ and $v(t)=S_{\mathcal{A}}(t)v_{0}$ with difference w=uvw=u-v. Then:
$\partial_{t}w$ $=\mathcal{A}w=-\nabla\cdot(\mathbf{v}w)+D\Delta w$ (23)
Taking the $L^{2}$ inner product with $w$:
$\frac{1}{2}\frac{d}{dt}\|w\|_{L^{2}}^{2}$ $=\langle w,-\nabla\cdot(\mathbf{v}w)\rangle+D\langle w,\Delta w\rangle$ (24)
For the advection term, using integration by parts and the boundedness of $\mathbf{v}$:
$|\langle w,-\nabla\cdot(\mathbf{v}w)\rangle|$ $\leq G_{1}\|\mathbf{v}\|_{L^{\infty}}\|w\|_{L^{2}}^{2}+\frac{D}{2}\|\Delta w\|_{L^{2}}^{2}$ (25)
where $G_{1}$ depends on the diffusion coefficient $D$ and the domain regularity.
For the diffusion term:
$D\langle w,\Delta w\rangle$ $=-D\|\nabla w\|_{L^{2}}^{2}\leq 0$ (26)
Thus:
$\frac{d}{dt}\|w\|_{L^{2}}^{2}$ $\leq 2G_{1}\|\mathbf{v}\|_{L^{\infty}}\|w\|_{L^{2}}^{2}$ (27)
By Gronwall’s inequality:
$\|w(\Delta t)\|_{L^{2}}^{2}$ $\leq e^{2G_{1}\|\mathbf{v}\|_{L^{\infty}}\Delta t}\|w(0)\|_{L^{2}}^{2}$ (28)
So for small $\Delta t$:
$\|S_{\mathcal{A}}(\Delta t)u_{0}-S_{\mathcal{A}}(\Delta t)v_{0}\|_{L^{2}}$ $\leq(1+L_{A}\Delta t)\|u_{0}-v_{0}\|_{L^{2}}$ (29)
where $L_{A}=2G_{1}\|\mathbf{v}\|_{L^{\infty}}$.
For the reaction equation $\partial_{t}u=r(u)$, consider two solutions $u(t)$ and $v(t)$ with difference w=uvw=u-v. Then:
$\frac{dw}{dt}$ $=r(u)-r(v)$ (30)
Using the Lipschitz condition $|r(u)-r(v)|\leq L_{r}|u-v|$ in Assumption 1:
$\left|\frac{d}{dt}\|w\|_{L^{2}}^{2}\right|$ $=2\left|\int w(r(u)-r(v))dx\right|\leq 2L_{r}\|w\|_{L^{2}}^{2}$ (31)
By Gronwall’s inequality:
$\|w(\Delta t)\|_{L^{2}}^{2}$ $\leq e^{2L_{r}\Delta t}\|w(0)\|_{L^{2}}^{2}$ (32)
So for small $\Delta t$:
$\|S_{\mathcal{R}}(\Delta t)u_{0}-S_{\mathcal{R}}(\Delta t)v_{0}\|_{L^{2}}$ $\leq(1+L_{R}\Delta t)\|u_{0}-v_{0}\|_{L^{2}}$ (33)
where $L_{R}=2L_{r}$.
Combining both estimates:
$\|S_{\mathcal{R}}(\Delta t)S_{\mathcal{A}}(\Delta t)\tilde{u}_{n}-S_{\mathcal{R}}(\Delta t)S_{\mathcal{A}}(\Delta t)u_{n}\|_{L^{2}}$
$\quad\leq(1+L_{R}\Delta t)\|S_{\mathcal{A}}(\Delta t)\tilde{u}_{n}-S_{\mathcal{A}}(\Delta t)u_{n}\|_{L^{2}}$
$\quad\leq(1+L_{R}\Delta t)(1+L_{A}\Delta t)\|\tilde{u}_{n}-u_{n}\|_{L^{2}}$
$\quad\leq(1+C_{1}\Delta t)\|\tilde{u}_{n}-u_{n}\|_{L^{2}}$ (34)
where $C_{1}=L_{A}+L_{R}+O(\Delta t)$. ∎
Lemma 2 (Splitting error estimate).
Under Assumption 1, the Lie splitting error satisfies:
$\mathbb{E}\,\|u^{\mathrm{Lie}}_{n+1}-u_{n+1}\|_{L^{2}}\leq C_{2}\Delta t^{2}$ (35)
where $C_{2}$ depends on the commutator $\mathcal{A},\mathcal{R}]$ and the solution regularity.
Proof.
Using the Baker-Campbell-Hausdorff formula [2, 5, 13] for operator splitting, we have the formal expansion:
$S_{\mathcal{R}}(\Delta t)\circ S_{\mathcal{A}}(\Delta t)-S_{\mathcal{A}+\mathcal{R}}(\Delta t)=\tfrac{\Delta t^{2}}{2}[\mathcal{R},\mathcal{A}]+\mathcal{O}(\Delta t^{3})$ (36)
where $\mathcal{R},\mathcal{A}]u=r^{\prime}(u)\mathcal{A}u-\mathcal{A}(r(u))$ is the commutator expression for the reaction and advection-diffusion operators.
For the specific operators $\mathcal{R}u=r(u)$ and $\mathcal{A}u=\nabla\cdot(\mathbf{v}u)+D\Delta u$, we compute the commutator explicitly:
$[\mathcal{R},\mathcal{A}]u$ $=r^{\prime}(u)\mathcal{A}u-\mathcal{A}(r(u))$
$=r^{\prime}(u)[\nabla(\mathbf{v}u)+D\Delta u]-\left[\nabla\cdot(\mathbf{v}r(u))+D\Delta r(u)\right]$
$=r^{\prime}(u)\nabla(\mathbf{v}u)+Dr^{\prime}(u)\Delta u$
$\quad-\left[r(u)\nabla\mathbf{v}+r^{\prime}(u)\mathbf{v}\cdot\nabla u+Dr^{\prime}(u)\Delta u+Dr^{\prime\prime}(u)|\nabla u|^{2}\right]$
$=r^{\prime}(u)\left[\nabla(\mathbf{v}u)-\mathbf{v}\cdot\nabla u\right]-r(u)\nabla\mathbf{v}-Dr^{\prime\prime}(u)|\nabla u|^{2}$
$=r^{\prime}(u)u\nabla\mathbf{v}-r(u)\nabla\cdot\mathbf{v}-Dr^{\prime\prime}(u)|\nabla u|^{2}$ (37)
Under Assumption 1, using Hölder’s inequality and Sobolev embedding theorem, we can obtain:
$\|[\mathcal{R},\mathcal{A}]u\|_{L^{2}}$ $\leq\|r^{\prime}(u)u\nabla\mathbf{v}\|_{L^{2}}+\|r(u)\nabla\mathbf{v}\|_{L^{2}}+D\|r^{\prime\prime}(u)|\nabla u|^{2}\|_{L^{2}}$
$\leq\|r^{\prime}(u)\|_{L^{\infty}}\|u\|_{L^{\infty}}\|\nabla\mathbf{v}\|_{L^{\infty}}|\Omega|^{1/2}+\|r(u)\|_{L^{\infty}}\|\nabla\mathbf{v}\|_{L^{\infty}}|\Omega|^{1/2}+D\|r^{\prime\prime}(u)\|_{L^{\infty}}\|\nabla u\|_{L^{4}}^{2}$
$\leq G_{2}(\|\mathbf{v}\|_{W^{1,\infty}},\|r\|_{C^{2}},\|u\|_{H^{s}})$ (38)
with the constant $G_{2}$ depending on $|\Omega|$, $D$, and the Sobolev embedding constants.
The local error bound of order $\Delta t^{2}$ follows immediately from the Baker-Campbell-Hausdorff formula. ∎
Lemma 3 (Density estimation error).
Under Assumptions 1 and 2, the density estimation error satisfies:
$\mathbb{E}\left[\|\widetilde{u}^{\mathrm{D}}_{n+1}-S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}\|_{L^{2}}\right]\leq C_{3}(N\Delta x^{d})^{-1/2}+C_{4}\Delta x$ (39)
where the constants $C_{3},C_{4}$ depend on the domain size $L$, dimension $d$, diffusion coefficient $D$, the total mass at $t_{n}$ $\widetilde{m}_{n}=\int_{\Omega}\widetilde{u}_{n}(\mathbf{x})d\mathbf{x}$, and the regularity parameters from Assumption 1.
Proof.
Let $\widetilde{M}_{n}=\int_{\Omega}\widetilde{u}_{n}(\mathbf{x})d\mathbf{x}$ be the computed total mass at time $t_{n}$, which contains error from previous steps. The density estimate in bin $j$ after the advection-diffusion step is:
$\hat{u}_{j}=\frac{\widetilde{m}_{n}\cdot K_{j}}{(\Delta x)^{d}}=\frac{\widetilde{M}_{n}}{N(\Delta x)^{d}}\cdot K_{j}$ (40)
where $\widetilde{m}_{n}=\widetilde{M}_{n}/N$ is the particle mass at time $t_{n}$ and $K_{j}$ is the number of particles in bin $j$.
The density estimation error consists of two components: statistical sampling error, and spatial discretization error.
1. Statistical sampling error: Let $p_{j}^{\mathrm{true}}=\int_{\mathrm{bin}_{j}}\frac{S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}(\mathbf{x})}{\widetilde{M}_{n}}d\mathbf{x}$ be the probability that a particle, starting from the initial distribution $\widetilde{u}_{n}$ and following the exact advection-diffusion dynamics, falls in bin $j$.
More precisely, consider particles $\mathbf{X}^{i}(t)$ satisfying the exact SDE:
$d\mathbf{X}^{i}(t)=\mathbf{v}(\mathbf{X}^{i}(t),t)dt+\sqrt{2D}d\mathbf{W}_{i}(t),\quad\text{for }t\in[t_{n},t_{n+1}]$ (41)
with initial distribution $\mathbf{X}^{i}(t_{n})\sim\widetilde{u}_{n}$. Then:
$p_{j}^{\mathrm{true}}=\mathbb{P}\left(\mathbf{X}^{i}(t_{n+1})\in\mathrm{bin}_{j}\mid\mathbf{X}^{i}(t_{n})\sim\widetilde{u}_{n}\right)$ (42)
The expected number of particles in bin $j$ is $\mathbb{E}[K_{j}]=Np_{j}^{\mathrm{true}}$, so the expected density is:
$\mathbb{E}[\hat{u}_{j}]=\frac{\widetilde{m}_{n}}{N(\Delta x)^{d}}\cdot\mathbb{E}[K_{j}]=\frac{\widetilde{m}_{n}}{N(\Delta x)^{d}}\cdot Np_{j}^{\mathrm{true}}=\frac{\widetilde{m}_{n}\cdot p_{j}^{\mathrm{true}}}{(\Delta x)^{d}}$ (43)
The variance of the density estimate is:
$\mathrm{Var}(\hat{u}_{j})=\left(\frac{\widetilde{m}_{n}}{N(\Delta x)^{d}}\right)^{2}\cdot\mathrm{Var}(K_{j})\leq\left(\frac{\widetilde{m}_{n}}{N(\Delta x)^{d}}\right)^{2}\cdot Np_{j}^{\mathrm{true}}(1-p_{j}^{\mathrm{true}})$ (44)
Using $p_{j}^{\mathrm{true}}\leq\frac{\|S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}\|_{L^{\infty}}(\Delta x)^{d}}{\widetilde{m}_{n}}$, we get:
$\mathrm{Var}(\hat{u}_{j})\leq\frac{\widetilde{m}_{n}\|S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}\|_{L^{\infty}}}{N(\Delta x)^{d}}$ (45)
Summing over all bins:
$\mathbb{E}\left[\sum_{j=1}^{M^{d}}\Delta x^{d}|\hat{u}_{j}-\mathbb{E}[\hat{u}_{j}]|^{2}\right]$ $=\sum_{j=1}^{M^{d}}\Delta x^{d}\cdot\mathrm{Var}(\hat{u}_{j})$
$\leq\sum_{j=1}^{M^{d}}\Delta x^{d}\cdot\frac{\widetilde{m}_{n}\|S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}\|_{L^{\infty}}}{N(\Delta x)^{d}}$
$\leq\frac{\widetilde{m}_{n}\|S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}\|_{L^{\infty}}(2L)^{d}}{N(\Delta x)^{d}}$ (46)
Taking square roots:
$\left(\mathbb{E}\left[\sum_{j=1}^{M^{d}}\Delta x^{d}|\hat{u}_{j}-\mathbb{E}[\hat{u}_{j}]|^{2}\right]\right)^{1/2}\leq\frac{C_{3}}{\sqrt{N}\Delta x^{d/2}}$ (47)
where $C_{3}=\sqrt{\widetilde{m}_{n}\|S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}\|_{L^{\infty}}(2L)^{d}}$.
2. Spatial discretization error: Let $u_{j}^{*}=S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}(\mathbf{x}_{j})$ be the true density at the $j^{th}$-bin center $\mathbf{x}_{j}$. The bias due to spatial discretization is:
$|\mathbb{E}[\hat{u}_{j}]-u_{j}^{*}|=\left|\frac{\widetilde{m}_{n}\cdot p_{j}^{\mathrm{true}}}{(\Delta x)^{d}}-S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}(\mathbf{x}_{j})\right|$ (48)
Note that:
$\frac{\widetilde{m}_{n}\cdot p_{j}^{\mathrm{true}}}{(\Delta x)^{d}}=\frac{1}{(\Delta x)^{d}}\int_{\mathrm{bin}_{j}}S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}(\mathbf{x})d\mathbf{x}$ (49)
By Taylor expansion and the smoothness of $S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}$ under Assumption 1:
$\left|\frac{1}{(\Delta x)^{d}}\int_{\mathrm{bin}_{j}}S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}(\mathbf{x})d\mathbf{x}-S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}(\mathbf{x}_{j})\right|\leq G_{3}\Delta x$ (50)
Thus, the $L^{2}$ spatial bias is:
$\left(\sum_{j=1}^{M^{d}}\Delta x^{d}|\mathbb{E}[\hat{u}_{j}]-u_{j}^{*}|^{2}\right)^{1/2}\leq C_{4}\Delta x$ (51)
where $C_{4}$ depends on the second derivatives of $S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}$ and the domain size.
Combining both error components Eqs.(47)(51):
$\mathbb{E}\left[\|\widetilde{u}^{\mathrm{D}}_{n+1}-S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}\|_{L^{2}}\right]\leq C_{3}(N\Delta x^{d})^{-1/2}+C_{4}\Delta x$ (52)
This completes the proof. ∎
Remark 2 (On the particle motion discretization error).
The particle motion discretization error from the Euler-Maruyama scheme, which would contribute a $\Delta t$ term to the density estimation error, is intentionally omitted from Lemma 3. This error arises from the discrepancy between the numerical particle positions $\mathbf{X}_{i}^{*}$ (following Euler-Maruyama) and the exact positions $\mathbf{X}_{i}^{\text{true}}(t_{n+1})$ (following the exact SDE dynamics).
However, this error can be treated as an accumulation term and is absorbed into the propagation error estimate in Lemma 1. Specifically, the $\Delta t$ coefficient in the propagation error bound $(1+C_{1}\Delta t)$ implicitly incorporates the effects of particle motion discretization.
Lemma 4 (Reaction integration error).
Under Assumption 1, the reaction integration error satisfies:
$\varepsilon_{\mathrm{react}}=\|\widetilde{u}^{\mathrm{R}}_{n+1}-S_{\mathcal{R}}(\Delta t)(\widetilde{u}^{\mathrm{D}}_{n+1})\|_{L^{2}}\leq\begin{cases}0,&\text{(exact integration)}\\ C_{5}\Delta t^{p},&\text{(numerical integration, order $p$)}\end{cases}$ (53)
where $C_{5}$ depends on the Lipschitz constant of $r(u)$ and the boundedness of higher derivatives under Assumption 1.
Proof.
Case 1: Exact integration For reactions with closed-form solutions (e.g., FKPP), the algorithm implements the exact solution operator:
$\widetilde{u}^{\mathrm{R}}_{n+1}=S_{\mathcal{R}}(\Delta t)(\widetilde{u}^{\mathrm{D}}_{n+1})\quad\Rightarrow\quad\varepsilon_{\mathrm{react}}=0$ (54)
Case 2: Numerical integration For general reactions, using a numerical method of order $p$ (e.g., implicit Euler: $p=1$, Crank-Nicolson: $p=2$) yields:
$\varepsilon_{\mathrm{react}}\leq C_{5}\Delta t^{p}.$ (55)
Lemma 5 (Resampling error).
Under Assumption 1, the resampling error satisfies:
$\mathbb{E}[\varepsilon_{\mathrm{resamp}}]=\mathbb{E}[\|\widetilde{u}_{n+1}-\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{2}}]\leq C_{6}(N\Delta x^{d})^{-1/2}$ (56)
where the constant $C_{6}$ depends on the domain size $L$, dimension $d$, and the total mass $M$.
Proof.
The resampling step converts the continuous density field $\widetilde{u}^{\mathrm{R}}_{n+1}$ back to a particle representation. This process introduces statistical sampling error:
The resampling step generates $N$ new particles according to the multinomial distribution with probabilities:
$p_{j}=\frac{\widetilde{u}^{\mathrm{R}}_{n+1,j}\cdot(\Delta x)^{d}}{M}$ (57)
where $M=\sum_{j=1}^{M^{d}}\widetilde{u}^{\mathrm{R}}_{n+1,j}\cdot(\Delta x)^{d}$ is the total mass.
Let $n_{j}$ be the number of particles assigned to bin $j$ with $\sum_{j=1}^{M^{d}}n_{j}=N$. The new density estimate after resampling is:
$\widetilde{u}_{n+1,j}=\frac{M\cdot n_{j}}{N(\Delta x)^{d}}$ (58)
The expected value is:
$\mathbb{E}[\widetilde{u}_{n+1,j}]=\frac{M}{N(\Delta x)^{d}}\cdot\mathbb{E}[n_{j}]=\frac{M}{N(\Delta x)^{d}}\cdot Np_{j}=\widetilde{u}^{\mathrm{R}}_{n+1,j}$ (59)
so the resampling is unbiased.
The variance is:
$\mathrm{Var}(\widetilde{u}_{n+1,j})=\left(\frac{M}{N(\Delta x)^{d}}\right)^{2}\cdot\mathrm{Var}(n_{j})=\left(\frac{M}{N(\Delta x)^{d}}\right)^{2}\cdot Np_{j}(1-p_{j})$ (60)
Using $p_{j}\leq\frac{\|\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{\infty}}(\Delta x)^{d}}{M}$, we get:
$\mathrm{Var}(\widetilde{u}_{n+1,j})\leq\frac{M\|\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{\infty}}}{N(\Delta x)^{d}}$ (61)
Summing over all bins:
$\mathbb{E}\left[\sum_{j=1}^{M^{d}}\Delta x^{d}|\widetilde{u}_{n+1,j}-\widetilde{u}^{\mathrm{R}}_{n+1,j}|^{2}\right]$ $\leq\sum_{j=1}^{M^{d}}\Delta x^{d}\cdot\frac{M\|\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{\infty}}}{N(\Delta x)^{d}}$
$\leq\frac{M\|\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{\infty}}(2L)^{d}}{N(\Delta x)^{d}}$
Taking square roots:
$\mathbb{E}[\|\widetilde{u}_{n+1}-\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{2}}]\leq\frac{C}{\sqrt{N}\Delta x^{d/2}}=C_{6}(N\Delta x^{d})^{-1/2}$ (62)
where $C_{6}=\sqrt{M\|\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{\infty}}(2L)^{d}}$.
Under Assumption 1, $M$ and $\|\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{\infty}}$ are uniformly bounded, ensuring $C_{6}$ is finite. ∎
Now, we are ready to prove Theorem 3.
Proof of Theorem 3. From Eqs.(15)(3), and all the lemmas above, we can obtain:
$\mathbb{E}\,\|\widetilde{u}_{n+1}-u_{n+1}\|_{L^{2}}$ $\leq\mathbb{E}[\varepsilon_{\mathrm{propag}}]+\mathbb{E}[\varepsilon_{\mathrm{dens}}]+\mathbb{E}[\varepsilon_{\mathrm{react}}]+\mathbb{E}[\varepsilon_{\mathrm{resamp}}]+\mathbb{E}[\varepsilon_{\mathrm{split}}]$
$\leq(1+C_{1}\Delta t)\mathbb{E}\,\|\tilde{u}_{n}-u_{n}\|_{L^{2}}+C_{3}(N\Delta x^{d})^{-1/2}+C_{4}\Delta x$
$\quad+C_{5}\Delta t^{p}+C_{6}(N\Delta x^{d})^{-1/2}+C_{2}\Delta t^{2}$ (63)
For the optimal case where we use exact reaction integration ($p\to\infty$) or second-order methods ($p=2$), we obtain:
$\mathbb{E}\,\|\widetilde{u}_{n+1}-u_{n+1}\|_{L^{2}}\leq(1+C_{1}\Delta t)\mathbb{E}\,\|\tilde{u}_{n}-u_{n}\|_{L^{2}}+C_{4}\Delta x+C_{7}\Delta t^{2}+C_{8}(N\Delta x^{d})^{-1/2}$ (64)
where $C_{7}=C_{2}+C_{4}$, $C_{8}=C_{3}+C_{6}$.
Since the initial condition is sampled exactly from $u_{0}$, we have $\varepsilon_{0}=0$. By the discrete Gronwall’s inequality, for $n\Delta t\leq T$:
$\mathbb{E}\,\|\tilde{u}_{n}-u_{n}\|_{L^{2}}$ $\leq\frac{e^{C_{1}T}-1}{C_{1}}\left(C_{7}(N\Delta x^{d}\Delta t)^{-1/2}+C_{8}\frac{\Delta x}{\Delta t}+C_{2}\Delta t\right)$
$\leq C_{T}\left(\Delta t+(N\Delta x^{d}\Delta t)^{-1/2}+\frac{\Delta x}{\Delta t}\right)$ (65)
where $C_{T}=\frac{e^{C_{1}T}-1}{C_{1}}\max(C_{7},C_{8},C_{2})$. This completes the proof of Theorem 3.
4 Numerical Experiments
4.1 1D Numerical Experiments
In this section, we present 1D numerical experiments to validate the robustness of the SGIP method for solving reaction-diffusion equations. We consider three different nonlinear reaction terms and compare the results with traditional FDM.
All experiments are conducted on the computational domain $\Omega=[-L,L]$ with $L=60$. The initial condition is given by a uniform distribution on the interval $[0,1]$:
$u_{0}(x)=\chi_{[0,1]}(x)=\begin{cases}1&\text{for }x\in[0,1]\\ 0&\text{otherwise}\end{cases}$ (66)
where $\chi_{[0,1]}$ denotes the characteristic function of the interval $[0,1]$. This discontinuous initial profile tests the method’s capability to handle sharp gradients and evolving interfaces.
For the reference FDM simulations, we employ a high-resolution discretization with spatial step $\delta x=10^{-3}$ and temporal step $\delta t=10^{-3}$ to ensure accurate reference solutions. For the SGIP, we use $N=10^{6}$ particles to approximate the density distribution, with the spatial domain discretized into $M=150$ bins for density estimation, corresponding to a bin width of $\Delta x=0.8$. The time step is set to $\Delta t=0.5$.
First, we consider the classical FKPP equation:
$u_{t}=u_{xx}+u(1-u).$ (67)
(a) t=5
(b) t=10
(c) t=15
(d) t=20
Figure 1: Comparison of FDM and SGIP method at $u_{t}=u_{xx}+u(1-u)$
Figure 1 displays the numerical solutions of the standard FKPP equation at different time instances. It can be observed that the SGIP method results show excellent agreement with FDM. As expected for the FKPP equation, the shape of the wave front remains unchanged during propagation, demonstrating the characteristic traveling wave structure.
Next, we consider a reaction-diffusion equation with cubic nonlinear reaction term:
$u_{t}=u_{xx}+u^{2}(1-u).$ (68)
(a) t=5
(b) t=10
(c) t=15
(d) t=20
Figure 2: Comparison of FDM and SGIP method at $u_{t}=u_{xx}+u^{2}(1-u)$
Figure 2 shows the numerical results for the case with cubic nonlinear reaction term. Notably, the cubic nonlinearity $u^{2}(1-u)$ leads to a visibly slower propagation speed compared to the standard FKPP case, while maintaining a stable wave profile. The SGIP method still accurately captures the wave front propagation behavior and maintains consistency with FDM results, demonstrating the robustness of the SGIP method when dealing with more complex nonlinear reaction terms.
Finally, we consider a reaction-diffusion equation with Arrhenius-type reaction term:
$u_{t}=u_{xx}+e^{-E/u}(1-u)$ (69)
where we choose $E=\frac{1}{2}$.
(a) t=5
(b) t=10
(c) t=15
(d) t=20
Figure 3: Comparison of FDM and SGIP method at $u_{t}=u_{xx}+e^{-1/2u}(1-u)$
Figure 3 presents the numerical results for the exponential reaction term case. This reaction term possesses singularity near $u=0$, posing additional challenges for numerical methods. The SGIP method demonstrates satisfactory performance in capturing the wave propagation under this singular nonlinearity. It is observed that the wave front propagates at a notably slower speed compared to the standard FKPP case, while maintaining a coherent wave profile. These results further validate the robustness of the SGIP in handling equations with complex nonlinearities.
To quantitatively compare the effects of different reaction terms on propagation speed, we plot the evolution of wave front position over time for all three cases in Fi gure 4. These numerical experiments comprehensively demonstrate the effectiveness and accuracy of the SGIP method for solving various types of reaction-diffusion equations.
(a) $u_{t}=u_{xx}+u(1-u)$
(b) $u_{t}=u_{xx}+u^{2}(1-u)$
(c) $u_{t}=u_{xx}+e^{-1/2u}(1-u)$
Figure 4: Propagation speed comparison for particle and FDM
4.2 2D Numerical Experiments
We now extend our investigation to 2D RAD systems to further validate the SGIP method’s capability in handling more complex spatial dynamics. The 2D experiments provide insights into pattern formation, wave propagation in multiple dimensions, and the interaction between diffusion and advection processes.
All two-dimensional simulations are conducted on the computational domain $\Omega=[-60,60]^{2}$. The initial condition is specified by sampling particle positions uniformly from the unit square: $\mathbf{X}_{0}^{i}\sim\text{Uniform}([0,1]^{2})$ for $i=1,\dots,N$, corresponding to an initial density $u_{0}(x,y)=\chi_{[0,1]^{2}}(x,y)$.
For the reference FDM simulations, we employ a high-resolution uniform grid with spatial discretization $\delta x=\delta y=10^{-2}$ and temporal step $\delta t=10^{-3}$ to ensure accurate reference solutions.
For the SGIP method, we use $N=3\times 10^{6}$ particles to approximate the density distribution. The spatial domain is discretized into $150^{2}$ uniform bins for density estimation, and the time step is set to $\Delta t=0.5$.
(a) t=5
(b) t=10
(c) t=15
(d) t=20
Figure 5: Comparison of FDM and SGIP method at $u_{t}=u_{xx}+u_{yy}+u(1-u)$
Figure 6: Propagation speed comparison for particle and FDM
Figure 5 demonstrates the evolution of the 2D FKPP equation with uniform initial conditions, showing contour lines at $u=0.1$, $0.5$, and $0.9$. The SGIP method successfully preserves the circular symmetry of the expanding wave front, maintaining rotational invariance in the contour morphology across all three concentration levels. Quantitative comparison in Figure 6 assesses the radial propagation speeds by tracking the intersection points of the $u=0.2$ contour with the x-axis for both SGIP and FDM. The excellent agreement between methods, evidenced by the linear position-time relationship, confirms the constant speed propagation characteristic of FKPP-type equations and validates the accuracy of our SGIP method in capturing two-dimensional front dynamics.
4.2.1 2D steady flow
We now investigate the interaction between reaction-diffusion dynamics and advection terms, which is particularly relevant in biological and environmental applications.
Shear Flow:
$u_{t}=D\Delta u+(\sin y,0)\cdot\nabla u+u(1-u).$ (70)
(a) SGIP
(b) FDM
Figure 7: SGIP vs. FDM on shear flow at t=20
Cellular Flow:
$u_{t}=D\Delta u+(-\sin x\cos y,\cos x\sin y)\cdot\nabla u+u(1-u).$ (71)
(a) SGIP
(b) FDM
Figure 8: SGIP vs. FDM on cellular flow
Cat’s Eye Flow:
$u_{t}=D\Delta u+[(-\sin x\cos y,\cos x\sin y)+\delta\cdot(\cos x\sin y,-\sin x\cos y)]\cdot\nabla u+u(1-u),$ (72)
here we set $\delta=2$.
(a) SGIP with D=0.5
(b) FDM with D=0.5
(c) SGIP with D=1
(d) FDM with D=1
Figure 9: SGIP vs. FDM on cat’s eye flow
Figure 7 demonstrates the effect of a sinusoidal shear flow on the reaction-diffusion dynamics, where the SGIP accurately captures the formation of characteristic shear-aligned patterns, showing how the flow stretches and orients the chemical waves along the shear direction. In Figure 8, the interaction with cellular flow patterns is shown, with the Lagrangian formulation naturally resolving the complex recirculation zones and their impact on front morphology. The cellular flow organizes the reaction fronts into approximately circular structures confined within individual flow cells, while inducing characteristic corrugations along the intercellular boundaries due to hyperbolic stagnation points and associated stretching fields. Figure 9 examines a more intricate cat eye flow combining cellular and shear components, where the SGIP successfully reproduces the competition between diffusion and advection across different diffusion parameter regimes ($D=0.5$ and $D=1$). The method captures the characteristic lobe structures formed by alternating elliptic islands and hyperbolic channels, with lower diffusivity ($D=0.5$) producing sharp lobe boundaries and higher diffusivity ($D=1$) enhancing cross-channel mixing and smoothing front morphology. The SGIP resolves how the flow superposition modulates front propagation through the competing recirculation and transport regions.
4.3 3D Numerical Experiments on ABC Flow
We now extend our investigation to 3D systems with the ABC flow, described by:
$u_{t}=D\Delta u+(A\sin z+C\cos y,B\sin x+A\cos z,C\sin y+B\cos x)\cdot\nabla u+u(1-u)$ (73)
where $A=1,B=\sqrt{\frac{2}{3}},C=\sqrt{\frac{1}{3}}$. Simulations are conducted on the domain $60,60]^{3}$ with initial condition $u_{0}(\mathbf{x})=\chi_{B(0,1)}(\mathbf{x})$, where $\chi_{B(0,1)}$ denotes the characteristic function of the unit ball centered at the origin.
The SGIP employs $5\times 10^{6}$ particles with $100^{3}$ spatial bins and time step $\Delta t=0.5$. In contrast, the FDM requires significantly finer discretization for numerical stability: time step $\delta t=10^{-3}$ and spatial grid of at least $300^{3}$ points, resulting in substantially higher computational cost.
(a) SGIP with projection at x=0 plane
(b) FDM with projection at x=0 plane
(c) SGIP with projection at y=0 plane
(d) FDM with projection at y=0 plane
(e) SGIP with projection at z=0 plane
(f) FDM with projection at z=0 plane
Figure 10: SGIP vs. FDM on ABC flow with $D=1$
(a) SGIP with projection at x=0 plane
(b) FDM with projection at x=0 plane
(c) SGIP with projection at y=0 plane
(d) FDM with projection at y=0 plane
(e) SGIP with projection at z=0 plane
(f) FDM with projection at z=0 plane
Figure 11: SGIP vs. FDM on ABC flow with $D=0.5$
(a) SGIP with projection at x=0 plane
(b) SGIP with projection at y=0 plane
(c) SGIP with projection at z=0 plane
Figure 12: SGIP on ABC flow with $D=0.1$
Figures 10 and 11 compare both methods on cross-sections through the coordinate planes for diffusion coefficients $D=1$ and $D=0.5$, showing excellent agreement. As $D$ decreases to 0.1 (Figure 12), advection and reaction dominate the dynamics, requiring even finer FDM discretization ($400^{3}$ grid points) that becomes computationally prohibitive due to excessive memory requirements on conventional workstations. The SGIP method, however, remains feasible and accurately captures the complex 3D flow structures across all diffusion regimes.
We further validate the SGIP using more complex nonlinear reaction terms (cubic, Arrhenius), maintaining the same computational parameters ($5\times 10^{6}$ particles, $100^{3}$ bins, $\Delta t=0.5$). However, since these new reaction terms result in markedly different reaction kinetics, the previously used unit ball initial condition would produce dynamics that are too confined and difficult to visualize effectively within the computational domain. To compensate for this and to generate clearly observable front propagation or pattern formation, the initial condition is modified to $u_{0}(\mathbf{x})=\chi_{B(0,5)}(\mathbf{x})$ on a larger support.
(a) SGIP with projection at x=0 plane
(b) FDM with projection at x=0 plane
(c) SGIP with projection at y=0 plane
(d) FDM with projection at y=0 plane
(e) SGIP with projection at z=0 plane
(f) FDM with projection at z=0 plane
Figure 13: SGIP vs. FDM on ABC Flow with Reaction Term: $u^{2}(1-u)$
(a) SGIP with projection at x=0 plane
(b) FDM with projection at x=0 plane
(c) SGIP with projection at y=0 plane
(d) FDM with projection at y=0 plane
(e) SGIP with projection at z=0 plane
(f) FDM with projection at z=0 plane
Figure 14: SGIP vs. FDM on ABC Flow with Reaction Term: $e^{-\frac{1}{2u}}(1-u)$
Figure 13 demonstrates the method’s performance with cubic reaction term $u^{2}(1-u)$, while Figure 14 shows results for the Arrhenius-type reaction $e^{-1/(2u)}(1-u)$. In both cases, the SGIP accurately captures the complex interplay between advection, diffusion, and nonlinear reaction dynamics, showing excellent agreement with FDM reference solutions.
The consistent performance across different reaction types and initial conditions demonstrates the robustness and versatility of the SGIP. Its Lagrangian framework naturally handles complex flow geometries and nonlinear reactions without requiring specialized discretization techniques. The method’s ability to maintain accuracy with relatively large time steps ($\Delta t=0.5$ versus FDM’s $\delta t=10^{-3}$) while avoiding the curse of dimensionality in 3D problems represents a significant computational advantage for complex ARD systems.
5 Conclusion
We have presented the SGIP method, a novel stochastic genetic interacting particle algorithm for solving reaction-diffusion-advection equations. The method’s effectiveness stems from its integrated design that combines three key components: operator splitting for process decoupling, stochastic particle dynamics for transport, and adaptive resampling for long-term stability.
The main theoretical contribution is a rigorous convergence analysis, where we prove that under appropriate assumptions, the method achieves consistent accuracy across various dimensionalities and problem configurations. This provides a solid mathematical foundation for the proposed methodology. Numerical experiments demonstrate SGIP’s capability in handling diverse challenges: from various reaction types (FKPP, cubic, Arrhenius) to complex flow configurations (shear, cellular, ABC flows) across 1D, 2D, and 3D domains. The Lagrangian formulation naturally handles sharp gradients and complex geometries without requiring specialized discretization techniques. Particularly in higher dimensions, the method offers computational advantages over traditional Eulerian approaches by avoiding the curse of dimensionality while maintaining accuracy with relatively large time steps.
The adaptive resampling strategy proves crucial for maintaining numerical stability and computational efficiency over long-time integration. Through intelligent particle redistribution guided by Eulerian density reconstruction, SGIP maintains an optimal balance between statistical accuracy and computational efficiency. Future work will focus on extending SGIP to multi-species systems, and implementing the fast adaptive algorithm discussed in Remark 1 to concentrate computational resources in active transition regions for enhanced computational efficiency in higher dimensions.
6 Acknowledgements
ZW was partly supported by NTU SUG-023162-00001, MOE AcRF Tier 1 Grant RG17/24. JX was partly supported by NSF grants DMS-2219904 and DMS-2309520, a Qualcomm Gift Award, and the Swedish Research Council grant no. 2021-06594 during his residence at the Institut Mittag-Leffler in Djursholm, Sweden, and the support during his stay at the E. Schrödinger Institute, Vienna, Austria, in the Fall of 2025. ZZ was supported by the National Natural Science Foundation of China (Projects 92470103 and 12171406), the Hong Kong RGC grant (Projects 17304324 and 17300325), the Seed Funding Programme for Basic Research (HKU), the Outstanding Young Researcher Award of HKU (2020–21), and the Seed Funding for Strategic Interdisciplinary Research Scheme 2021/22 (HKU).